149
edits
No edit summary |
No edit summary |
||
--[[User:Zjsimmons|Zjsimmons]] ([[User talk:Zjsimmons|talk]]) 16:11, 20 May 2014 (CDT)
Trying to get fast feedback to laser diode in place to see if it helps with cavity locking. So far it hasn't worked/i haven't been able to get good slow feedback locking. Difficult to untangle the different knobs we have control over. Issues/brainstorming to consider:
# Tune up cavity peaks (IR) with crystal temp de-tuned and then zero in on correct temp? My suspicion is that adjacent peaks from different orders are messing up the error signal which is why it maybe isn't working well. If we tune up the IR, maybe this would be better to look at than trying to look at the green.
|